### v2

# Gerhard Jäger

Abstract
Computational phylogenetics has become an established tool in historical linguistics, with many language families now analyzed using likelihood-based inference. However, standard approaches rely on expert-annotated cognate sets, which are sparse, labor-intensive to produce, and limited to individual language families. This paper explores alternatives by comparing the established method to two fully automated methods that extract phylogenetic signal directly from lexical data. One uses automatic cognate clustering with unigram/concept features; the other applies multiple sequence alignment (MSA) derived from a pair-hidden Markov model. Both are evaluated against expert classifications from Glottolog and typological data from Grambank. Also, the intrinsic strengths of the phylogenetic signal in the characters are compared. Results show that MSA-based inference yields trees more consistent with linguistic classifications, better predicts typological variation, and provides a clearer phylogenetic signal, suggesting it as a promising, scalable alternative to traditional cognate-based methods. This opens new avenues for global-scale language phylogenies beyond expert annotation bottlenecks.
Beyond cognacythanks: This is an updated version of the paper published in the Proceedings of SIGTYP 2025, correcting a sign error in the pHMM training code and two bugs in the T-Coffee MSA construction. The qualitative conclusions are unchanged.
Gerhard Jäger University of Tübingen, Germany gerhard.jaeger@uni-tuebingen.de March 27, 2026thanks: Published in Proceedings of The 7th Workshop on Research in Computational Linguistic Typology and Multilingual NLP
1 Introduction
Originally developed in computational biology, quantitative methods for phylogenetic reconstruction using likelihood-based inference frameworks have now gained widespread acceptance in comparative linguistics. This is evident from the growing number of computational phylogenies proposed for some of the world's largest language families, including Dravidian (Kolipakam et al., 2018), Sino-Tibetan (Sagart et al., 2019), and Indo-European (Heggarty et al., 2023). Moreover, fully automated approaches — where even cognate identification is performed algorithmically — have demonstrated a surprising degree of robustness (Rama et al., 2018). In contrast to the pre-computational era of historical linguistics, where such detailed reconstructions were rare, the generation of fully resolved phylogenies with branch lengths and, in some cases, estimated divergence dates has now become a standard practice in studies of language evolution.
Despite the increasing recognition of computational language phylogenies as a useful addition to the comparative linguistics toolkit, skepticism remains prevalent. A key concern raised by critics is that phylogenetic analyses are often based on cognate sets—groups of historically related words—extracted from semantically aligned word lists. Since these cognate sets are based on expert annotations, they are sparse, labor-intensive to acquire, and raise concerns regarding replicability.
Another limitation of phylogenetic inference based on cognate classes is that it is by definition constrained to individual language families. There is legitimate interest in automatically inferred trees spanning larger collections of languages, perhaps from the entire world. Such trees provide information about the strength of evidence for putative macro-families (Jäger, 2015; Akavarapu and Bhattacharya, 2024). Furthermore, they are useful for downstream tasks such as the statistical modeling of global language evolution (Bentz et al., 2018; Bouckaert et al., 2022).
The literature contains several proposed workflows for extracting character matrices from word lists without cognate annotations, which can then used as input for likelihood-based phylogenetic inference. This paper presents a comparison of cognate-based phylogenetic inference with two such proposals, the one by Jäger (2018) and the one by Akavarapu and Bhattacharya (2024). These methods are evaluated in three ways: (1) by comparing the inferred phylogenies with the Glottolog expert classification (Hammarström et al., 2024), (2) how well the inferred phylogenies fit to the typological features from (Skirgård et al., 2023), and (3) an estimation of the strength of the phylogenetic signal in the data, which is inferred with the software PyPythia (Haag et al., 2022).
2 Materials and Methods
2.1 Materials
Word lists were obtained from Lexibank [^1] https://github.com/lexibank; List et al. 2023. These datasets contain lexical entries, including the language they belong to, their meaning, form in IPA transcription, and often a manual cognate annotation. The datasets are curated by the Lexibank community and are available in a standardized format, which makes them suitable for computational analyses.
In a first step, 135 Lexibank dataset were selected. In total, this amounts to 2,486,845 lexical entries from 6,845 languages (identified by glottocodes).
For the purpose of evaluation, typological features were obtained from Grambank [^2] https://github.com/grambank/grambank. This results in 355,097 binary entries from 2,467 languages and 195 typological features.
A subset of Lexibank data was selected according to the following criteria:
  • The entry comes from a language with a Glottocode that is present in the Grambank data.
  • The entry has an entry for its meaning (Concepticon_Gloss) and a manual cognate annotation (Cognateset_ID).
  • The meaning comes from the 110 concepts with the largest coverage.
This leaves 113,671 entries from 928 languages. For further processing, the IPA transcriptions were converted to the ASJP alphabet using the python package lingpy (List and Forkel, 2024).
Constraining the Grambank data to these 928 languages leaves 138,878 binary data points from all 195 features.
The gold standard tree was obtained from Glottolog. [^3] https://zenodo.org/records/10804582/files/glottolog/glottolog-cldf-v5.0.zip
2.2 Methods
The overall workflow consists of the following steps:
  1. 1.
    Phylogenetic inference
    1. (a)
      Generate a binary character matrix from the Lexibank data.
    2. (b)
      Infer a phylogenetic tree from this character matrix.
  2. 2.
    Evaluation
    1. (a)
      Compare the inferred phylogenetic tree with the Glottolog expert classification.
    2. (b)
      Compare the inferred phylogenetic tree with the Grambank typological features.
    3. (c)
      Assess the strength of the phylogenetic signal in the data.
Three different methods were used to generate a binary character matrix: (1) binarized expert-annotated cognate classes, (2) a combination of automatic cognate clustering and unigram/concept features as described in Jäger (2018), and (3) a variant of the method developed by Akavarapu and Bhattacharya (2024) using multiple sequence alignment.
2.2.1 Expert-annotated cognate classes (cc)
Here we use the method introduced by Ringe et al. (2002) and Gray and Atkinson (2003). Each cognate class is treated as a character. A language is coded as 1 if it has a cognate in the class, 0 if it has a different cognate class for the same concept, and missing if it has no cognate for the concept. This results in a matrix with 928 rows and 25,913 columns.
Since each cognate class is, by definition, confined to a single language family, this character matrix contains no signal beyond the family level.
In the tables and figures below, this method is referred to as cc (for cognate classes).
2.2.2 Automatic cognate clustering and unigram/concept features (PMI)
The workflow proposed by Jäger (2018) was replicated. This approach uses two types of characters.
  • Binarized cognate classes obtained via automatic cognate clustering. This involves (1) supervised training of a Support Vector Machine classifier which takes a pair of words and predicts the labels 1 (cognate) or 0 (non-cognate), using manual cognate classification for supervision, (2) creating a distance matrix for all entries for a given concept from the 100 concepts defined above, and (3) clustering the distance matrix using the label propagation algorithm (Raghavan et al., 2007).
  • Unigram/concept characters. For each combination of a concept $c$ and an ASJP sound class $s$, a language is coded as 1 if it has a word for concept $c$ that contains sound class $s$, missing if it has no word for concept $c$, and 0 otherwise.
This resulted in a matrix with 928 rows and 41,013 columns.
Since the pointwise mutual information between sound classes plays an essential role in this workflow, the method is referred to as PMI.
2.2.3 Multiple sequence alignment (MSA)
The method by Akavarapu and Bhattacharya (2024) was used as starting point, but the present approach differs in various aspects. The method is based on the following steps:
In a first step, pairwise distances between languages in the full lexibank dataset were computed using the Levenshtein distance on the ASJP transcriptions and aggregating according to the method described in (Jäger, 2018). Language pairs with a distance below $0.7$ were considered as probably related, using the same heuristics as Jäger (2018). There are 172,681 such language pairs. All word pairs from such a language pair sharing their meaning are treated as potential cognates. There are 90,565,486 such word pairs. An equal number of random word pairs were sampled as probable non-cognates. Potential cognates were assigned the label 1 and probable non-cognate the label 0.
In a second step, a classifier was trained on the potential cognates and non-cognates. The classifier consists of a pair-Hidden Markov Model (pHMM) (Durbin et al., 1998) and a logistic-regression layer. The classifier was trained for one epoch using the Adam optimizer. The resulting parameters of the pHMM were used in the next step.
A pHMM defines a probability distribution over pairs of aligned sequences of sound classes. This involves (1) emission probabilities for all pairs of sound classes that are matched in the alignment, (2) emission probabilities for individual sound classes if they are aligned with a gap, and (3) transition probabilities between the hidden states match, gap in string 1, gap in string 2, and final state.
It is instructive to inspect the emission probabilities in the trained model. In Table 1 the ten sound classes with the highest probability of being matched with /p/ are shown for illustration, together with their log-probabilities. This ranking is in good agreement with linguistic intuitions about potential sound correspondences.
Sound class Log-probability
p 2.23-2.23
f 11.06-11.06
b 20.25-20.25
v 20.52-20.52
6 26.72-26.72
2 28.58-28.58
3 28.69-28.69
c 29.76-29.76
L 30.33-30.33
u 30.41-30.41
Table 1: The ten sound classes with the highest probability of being matched with /p/ in the trained pHMM, along with their log-probabilities.
Sound class Log-probability
2 -0.19
1 -2.09
3 -3.06
q -6.12
E -6.47
5 -7.05
x -7.73
i -9.38
7 -9.79
G -10.64
Table 2: The ten sound classes with the highest probability of being matched with a gap in the trained pHMM, along with their log-probabilities. In List's version of ASJP, numerals 1–6 encode tone contours (level, rising, falling, etc.) and 7 encodes the glottal stop ().
A high probability here is to be interpreted as a high likelihood that instances of these sound classes participate either in insertion or deletion.
The trained pHMM assigns a probability to each pair of aligned sequences. Via the forward algorithm, the probability of a pair of sequences is computed as the sum of the probabilities of all possible alignments between these sequences.
Following Durbin et al. (1998), a null-model was trained additionally that assigns individual probabilities to both sequences, disregarding the order of sound classes. The log-odds ratio of a pair of words of being generated by the pHMM vs. the null model can interpreted as a measure of the similarity of the two words.
To illustrate this, a collection of ten words were chosen at random from the dataset which all have an edit distance of 1 to the word baba, and their log-odds ratios with respect to baba were computed. The results are shown in Table 3.
word log-odds
bIba -49.37
babae -49.44
babau -52.13
eaba -62.23
bawa -62.79
zaba -66.39
raba -76.10
xaba -77.01
maba -79.03
naba -90.25
Table 3: Ten randomly chosen words with an edit distance of 1 from baba, alongside with the predicted log-odds to baba.
This ranking illustrates that the log-odds predicted by the trained pHMM are consistent with linguistic intuitions about potential cognacy.
In a third step, the trained pHMM was used in combination with the Viterbi algorithm to obtain pairwise sequence alignments for all synonymous word pairs from different languages within the smaller dataset of 928 languages and 110 concepts.
In a fourth step, the pairwise sequence alignments were aggregated to a multiple sequence alignment (MSA) using the T-Coffee algorithm (Notredame et al., 2000).
Note that all reflexes of a given concept are aligned within a single MSA, regardless of cognacy. Such an MSA implicitly contains information both about cognacy and about sound correspondences.
An example (for a much smaller dataset) is shown in Table 4 for illustration. These are the reflexes of the concept louse from the Tungusic languages in the dataset. [^4] The data are taken from https://zenodo.org/records/13163376, which is based on (Oskolskaya et al., 2021).
Language Cognateset_ID 1 2 3 4 5 6 7 8
Even 16_lousen-37 k - u - m - k e
Kilen 16_lousen-37 q h u - m I k I
Negidal 16_lousen-37 k - u - m - k I
Oroch 16_lousen-37 k - u - m - - I
Udihe 16_lousen-37 k - u - m u x I
Nanai 16_lousen-38 t - i k t - - I
Orok 16_lousen-38 t - i k t - - I
Ulch 16_lousen-38 t - i q t - - I
Language sound class k q
Even 0 0 0
Kilen 0 0 0
Negidal 0 0 0
Oroch 0 0 0
Udihe 0 0 0
Nanai 1 1 0
Orok 1 1 0
Ulch 1 0 1
Table 4: Example of a multiple sequence alignment. Alignment cells are shaded to indicate different cognate sets. (left) Binarized version of column 4. (right)
As can be seen from this example, the MSA contains information about cognacy, but also about sound correspondences. For example, a t in the first column is a proxy for the cognate class 16_lousen-38. The sound classes k and q, on the other hand, both correspond to the cognate class 16_lousen-37, and they additionally reflect a sound change. In column 4, however, the cognate class 16_lousen-38 is split into two sound classes, k and q, reflecting a sound change. The presence of a sound class, as opposed to a gap, is a proxy of that cognate class. Put differently, binary characters corresponding to a gap are flipped by switching 0s and 1s.
In a fifth step, the MSA was converted to a binary matrix. Two binarization methods were used simultaneously. For a given column in an MSA, a character was created for the presence of a sound class. For column 4 in Table 4, e.g., this character has value $1$ for Nanai, Orok and Ulch, and 0 for the other languages. Additionally, for each sound class type in a column, a different character was created. In the example, there are two such characters, one for k and one for q. The first has value $1$ for Nanai and Orok and 0 otherwise, while the second has value $1$ for Ulch and 0 otherwise. Languages for which the data do not contain a reflex for a given concept are coded as missing for all relevant characters. If a language has multiple reflexes for a given concept, the maximum value is chosen.
Applying this workflow to all concepts and concatenating the resulting matrices yields the final character matrix 928 rows and 54,355 columns.
As mentioned above, this workflow builds on the method by Akavarapu and Bhattacharya (2024), but differs in various aspects. The mentioned work (1) uses Dolgopolsky sound classes instead of ASJP, (2) finds the MSA using CLUSTALW2 (Larkin et al., 2007) instead of T-Coffee, and (3) omits the binarization steps, working with a multi-state model of evolution for phylogenetic inference.
In the tables and figures this method is referred to as MSA.
2.2.4 Phylogenetic inference
We performed phylogenetic inference using raxml-ng (Kozlov et al., 2019), which implements maximum-likelihood estimation. The BIN+G model (binary substitution model with gamma-distributed rates) was used for all analyses. This means that gain rates and loss rates can be different, and that the mutation rates of the different characters can differ but are drawn from the same gamma distribution. The parameters of this distribution are estimated from the data.
Using the standard settings of raxml-ng, 20 maximum likelihood tree searches were performed, ten of them starting from random trees and ten from maximum-parsimony trees. The tree with the highest likelihood was chosen as the final result.
Method GQD (Glottolog) AIC (Grambank) difficulty
Cognate classes $0.188$ $105.340$ $0.59$
PMI $0.062$ $104.903$ $0.63$
MSA 0.054 104,787 0.21
Table 5: Evaluation of the full dataset. GQD = Generalized Quartet Distance (lower is better; ranges from 0 for perfect fit to $0.67$ for chance level); AIC = Akaike Information Criterion for typological model fit (lower is better; absolute values are not interpretable in isolation but differences are meaningful); difficulty = phylogenetic difficulty estimated by PyPythia (lower is better; ranges from 0 for strong phylogenetic signal to 1 for absent signal).
Method $\mu$ GQD $\sigma$ GQD $\mu$ AIC $\sigma$ AIC $\mu$ difficulty $\sigma$ difficulty
Cognate classes $0.227$ $0.077$ $151$ $115$ $0.486$ $0.030$
PMI $0.095$ $0.030$ 28-28 $69$ $0.326$ $0.032$
MSA 0.053 $0.013$ -129 $72$ 0.201 $0.050$
Table 6: Evaluation of the 100 random samples ($\mu$: sample mean; $\sigma$: sample standard deviation).
Method $\mu$ GQD $\sigma$ GQD $\mu$ AIC $\sigma$ AIC $\mu$ difficulty $\sigma$ difficulty
Cognate classes $0.223$ $0.130$ 1.73-1.73 $17.01$ $0.401$ $0.164$
PMI $0.221$ $0.109$ $3.42$ $20.43$ $0.280$ $0.187$
MSA 0.215 $0.110$ -4.45 $12.68$ 0.226 $0.110$
Table 7: Evaluation of the 14 largest language families ($\mu$: sample mean; $\sigma$: sample standard deviation).
Figure 1: Left panel: Comparison of methods across three evaluation metrics for the 100 random samples. The boxplots show distribution per method, while the overlaid points represent individual samples. Right panel: Comparison of methods across three evaluation metrics for the 14 largest language families. The boxplots show distribution per method, while the overlaid points represent individual samples.
2.2.5 Evaluation
Evaluation was conducted on three types of datasets:
  • the full dataset of 928 languages,
  • 100 samples of 100 languages each, which are drawn at random without replacement from the full dataset, and
  • a collection of 14 language families, each with at least 10 languages.
For each of these groups of datasets, the following evaluations were performed:
Comparison with Glottolog
The Glottolog classification of the languages in a dataset can be represented as a phylogenetic tree with polytomies, i.e., with nodes containing more than two daughters. This Glottolog tree serves as gold standard. To assess the degree of agreement between the gold standard and the inferred phylogenies, the generalized quartet distance (GQD) was deployed, as first proposed by Pompei et al. (2011). This distance is defined as the fraction of quartets (i.e., sets of four languages) that are (a) resolved in both trees, and (b) resolved differently in the two trees. The GQD ranges from 0 (perfect agreement) to $0.67$ (chance level). The GQD was computed using the software QDist, which can be obtained from https://birc.au.dk/software/qdist/.
Fit with Grambank
The hypothesis is assumed that the values of the Grambank features evolve along a phylogeny in the same way as the lexical characters described earlier in this section. The degree of fit of the inferred phylogenies with the Grambank features was assessed by (1) using the inferred phylogeny and estimating the branch lengths, mutation rates and rate heterogeneity via Maximum Likelihood, and (2) computing the Akaike Information Criterion (AIC). A lower AIC value indicates a better fit.
ML inference and AIC computation were also performed with raxml-ng.
For the groups of random samples and of language families, AIC values were normalized to mean 0 to facilitate comparison.
Phylogenetic difficulty
The strength of the phylogenetic signal in the data was assessed using the software PyPythia (Haag et al., 2022). The authors define a measure of signal strength that uses 100 maximum likelihood tree searches and quantifies the degree of agreement between the inferred trees. The software PyPythia implements a machine learning algorithm that predicts this difficulty from various properties of the character matrix, such as entropy and sites-over-taxa ratio, and maximum-parsimony tree inference, with high precision and comparatively low computational cost. The measure ranges from 0 (little difficulty, i.e., strongest signal) to 1 (very difficult, i.e., no signal).
3 Results
The evaluation results for the entire dataset are shown in Table 5. Table 6 shows the aggregated results for the 100 random samples. They are visualized in the left panel of Figure 1.
The aggregated evaluation results for the 14 largest language families are shown in Table 7. They are visualized in the right panel of Figure 1. The results for the individual families are given in Table 8.
Family Metric $c$ pmi msa
Afro-Asiatic GQD 0.455 0.195 0.212
PhyDiff 0.420 0.260 0.200
AIC 12.635 8.841 -29.519
Arawakan GQD 0.280 0.346 0.317
PhyDiff 0.250 0.110 0.120
AIC 0.868 -0.845 1.088
Atlantic-Congo GQD 0.150 0.220 0.242
PhyDiff 0.680 0.670 0.300
AIC -34.076 15.572 5.987
Austroasiatic GQD 0.052 0.093 0.095
PhyDiff 0.280 0.310 0.270
AIC 11.338 -16.115 -3.022
Austronesian GQD 0.161 0.200 0.160
PhyDiff 0.670 0.600 0.370
AIC -31.419 65.299 -28.045
Chibchan GQD 0.118 0.332 0.360
PhyDiff 0.350 0.360 0.290
AIC -0.168 -3.895 3.884
Dravidian GQD 0.312 0.242 0.261
PhyDiff 0.390 0.170 0.200
AIC 1.196 -0.862 -0.343
Indo-European GQD 0.031 0.014 0.016
PhyDiff 0.320 0.210 0.260
AIC 1.065 -19.079 8.237
Pama-Nyungan GQD 0.178 0.359 0.251
PhyDiff 0.540 0.420 0.450
AIC -17.375 13.601 1.963
Sino-Tibetan GQD 0.230 0.318 0.245
PhyDiff 0.590 0.300 0.260
AIC 32.455 -13.239 -23.957
Tucanoan GQD 0.421 0.274 0.400
PhyDiff 0.270 0.030 0.040
AIC -1.364 0.758 0.564
Tupian GQD 0.353 0.294 0.255
PhyDiff 0.390 0.200 0.180
AIC -0.187 -0.280 -0.001
Turkic GQD 0.249 0.117 0.117
PhyDiff 0.350 0.230 0.120
AIC -1.266 0.647 0.612
Uto-Aztecan GQD 0.126 0.084 0.076
PhyDiff 0.120 0.050 0.110
AIC 2.098 -2.485 0.309
Table 8: Evaluation of the 14 largest language families. The best value for each family is highlighted in bold.
When focusing on phylogenetic inference at the level of individual families, we find a considerable variation between families. This applies both to the numerical evaluation results and the relative ranking of the three methods considered here. The MSA method tends to produce the lowest phylogenetic difficulty, while there is no discernible trend regarding the fit to Glottolog and to Grambank.
This picture changes considerably when we focus on datasets covering languages from many different families. Here, the MSA method consistently outperforms the other two methods. This is particularly evident in the comparison with Glottolog, where the MSA method yields the lowest GQD values. The MSA method also leads to the lowest AIC values, indicating a better fit to the Grambank typological features. The phylogenetic difficulty is also lowest for the MSA method.
4 Discussion
These findings suggest that the MSA method is a promising alternative to traditional cognate-based methods. It is competitive with the more labor-intensive method based on manual cognate annotations, as well as the method using automatically detected cognate classifications, when considering individual language families. For global datasets, the MSA method clearly outperforms the other two methods. This is particularly evident in the comparison with Glottolog, where the MSA method yields the lowest GQD values. The MSA method also tends to produce the lowest AIC values, indicating a better fit to the Grambank typological features. The phylogenetic difficulty is also lowest for the MSA method.
Limitations
The two evaluation methods that quantify the fit of the inferred trees to empirical data only assess the quality of the inferred tree topologies. Future work will need to address the question how well the inferred branch lengths and divergence dates correspond to the true values. This is a challenging task, as the true values are unknown. It is expected that the usefulness for downstream tasks is a suitable proxy.
Data and Code Availability
The code used in this study is available at https://codeberg.org/profgerhard/sigtyp2025_code/.
Acknowledgments
This research was supported by the DFG Centre for Advanced Studies in the Humanities Words, Bones, Genes, Tools (DFG-KFG 2237) and by the European Research Council (ERC) under the European Union's Horizon 2020 research and innovation programme (Grant agreement 834050).